OpenStudy (warriorz13):

A race car driver lost 8 seconds off the leader for each of the 3 pit stops the driver made. At each of the next 2 pit stops the driver gained 10 seconds on the leader. Which represents the amount of time the driver is off the leader? –4 seconds –2 seconds 20 seconds 54 seconds

OpenStudy (anonymous):

-8*3=-24 -24+10(2)=-4 so A

OpenStudy (warriorz13):

TY

OpenStudy (anonymous):

You are very welcome